20160105 » History » Version 14

Version 13 (Boyd, Steven, 05 January 2016 17:30) → Version 14/23 (Boyd, Steven, 11 January 2016 09:03)

h1. Commissioning and Ops Scheduling meeting : 12/1/2016 5/1/2016

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

h2. Agenda:

* Review of tasks and activities identified at last meeting
** Realignment test of the SSD
Q4 flush - take magnet to high current and see if it trips.
** DC to contact Health Physics about calibration Temperature readout investigation/swap of neutron monitor whilst we have access Q4 / Q5 sensors (Henry)
** Simulation of effect of air Network reconfiguration for C&M : Document by Pierrick on changes are needed in the drift space between tracker and absorber network configuration so that system owners
understand the impact.
** Talk to John about conventional magnet cooling system flush
** Henry/Pierrick to come up with a design for the Q456 drip trays + finish Q789 drip trays
** Design for safe access to top of PRY / Andy
* Upcoming activities : chaired by DC
** [[WorkPlanJanuary2016|To-do list]] for Maintenance shutdown (January 2016)
** "Operations calendar":
** Looking ahead
*** ISIS run-up starts Feb 1
*** Next user cycle : Feb 19th - March 25th
*** Expect to take data from the last week of February (no channel magnets on, no DS). Initially plan to take data with
an empty absorber, but would like to change to LiH in early March.
* Date of next meeting : January 12th

h2. Minutes

Present: Holge, Craig, Pierrick, Mark T, Andy, John G, Victoria B, Henry, Ed, Colin, Josef, Steve, Ken

* Review of tasks and activities:
** Q4 was flushed and operated and maximum current. It didn't trip (which doesn't necessarily mean that the trip problem was fixed, just that we haven't harmed Q4).
** Pierrick checked the temperature readout for Q5/6 which appears to be swapped. He believes that the problem is in the addressing, and
this can be fixed in software. He's implemented a fix, but hasn't tested it yet.
** Pierrick has written up a document for the changes to the network for C&M but has not distributed it yet. He thinks that the work will
take place from Jan 15th - Jan 31st.
** John has already started flushing the other conventional magnets but has run into problems with air blocks in the line. He thinks he has
a work around for this which he'll look at this week.
** Henry has come up with a design for the conventional magnet drip trays and will talk to John about it on Thursday.
** Andy and John are on discussions about a way to make the top of the PRY safe. At present the idea is to install some mesh to
stop people falling into the PRY from above with a N-S clip-on bridge, but this needs some development. Hope to have this in place by
the end of the month. Josef has used a Scottish company which manufactures bespoke bridges like this and will talk to Andy.

* Activities this month.
** The operations calendar has been updated with activities that are planned for this month.
** There was discussion concerning whether SSD can be manually realigned before the Feb data-taking or between the Feb
and April user cycles. It was clear that we wouldn't decide in this meeting
*** *Action Item* : Colin will organise a meeting for Thursday morning to consider this issue.
** *Action Item* : DC to contact health physics about calibration of the neutron monitor whilst we have access (i.e. before the end of the month)
** *Action Item* : A simulation is needed to work out the effects of having half an atmosphere of air in the drift space ( & for the
matter the same pressure of He or Nitrogen in the LiH space) with the bellows removed.